OpenStudy (anonymous):

prime factorization of 8x10^18

OpenStudy (thomas5267):

What are the prime factors of 8 and 10?

OpenStudy (anonymous):

No like 8 * 10\[10^{18}\]

OpenStudy (anonymous):

8 times 10 to the 18 power

OpenStudy (thomas5267):

\[ \begin{align*} &8\times10^{18}\\ =&2^3\times(2\times5)^{18}\\ =&2^3\times2^{18}\times5^{18}\\ =&2^{21}\times5^{18} \end{align*} \]

OpenStudy (anonymous):

Thanks(: